# Bimanual Deformable Bag Manipulation Using a Structure-of-Interest Based Latent Dynamics Model

### v1

@@ IV-A Global Particle Sampling from Raw Observation
In this work, we begin by sampling representative particles from raw visual inputs, i.e., RGB-D images, to facilitate the training of Graph Neural Network (GNN)-based dynamics model, depicted in Fig. 4(a). The challenge lies in extracting meaningful particle representations of the Structure of Interest (SOI) from visual data that is significantly occluded. To this end, we introduce a comprehensive global particle sampling methodology, as illustrated in Fig. 5.
Preprocessing: The initial phase of our methodology involves refining the raw point cloud data obtained from RealSense D435 cameras. With appropriate calibration, we can transform the raw RGBD images into point cloud data with the cameras’ intrinsic and extrinsic parameters. Through meticulous calibration, we can convert raw RGB-D images into point cloud data, utilizing the intrinsic and extrinsic parameters of the cameras. To enhance data quality, a statistical outlier removal filter is applied to strip away noise, paving the way for accurate analysis. Focusing on the bag’s opening rim, which is crafted from green fabric, we employ a color-based segmentation algorithm. A specific color threshold within the HSV color space is set to discern green hues, effectively isolating the rim from the rest of the bag. Concurrently, we identify and omit the points marking the bag’s handle areas where robots will interact with the fabric using a separate HSV threshold. To ensure the rim’s point cloud is distinctly segregated from other bag components, Euclidean clustering is performed as a concluding step, particularly if handle detection leaves any connected segments. This meticulous process results in a purified dataset, optimal for the rim’s geometric representation.
Surface Reconstruction: To reconstruct the green rim’s surface from its point cloud, we leverage Poisson reconstruction, a preferred technique for its proficiency in managing occlusions by interpolating over incomplete data to form a smooth, continuous surface. Traditional methods like ball pivoting and alpha shapes underperform in occluded scenarios. Before reconstruction, we compute point normal via local principal axis analysis to guide the Poisson algorithm. The resulting surface is further polished with a uniform resampling algorithm, ensuring mesh uniformity and feature preservation for subsequent modeling and manipulation tasks.
Refinement with Topological Prior: In this step, we ensure that the reconstructed surface conforms to the anticipated topological structure of the bag’s opening rim through a two-fold approach: 1) Topological Analysis: We incorporate topological priors by recognizing that the rim should form a contiguous loop. This knowledge is employed to scrutinize the mesh, detecting any topological discrepancies. 2) Corrective Actions: Upon identifying any topological deviations, we undertake corrective measures by cutting and remeshing to realign the mesh with a loop topology that accurately represents the bag’s rim.
@@
Resampling: In the final stage of our process, our objective is to generate a high-quality particle set that accurately represents the geometry of the rim. To achieve this, we initially implement voxel grid down-sampling to attain a uniform distribution of vertices while concurrently eliminating statistical outliers. Subsequently, we employ the farthest point sampling technique to selectively reduce the point cloud associated with the Structure of Interest (SOI) to a manageable quantity conducive to Graph Neural Network (GNN) training. The culmination of this process, resulting in a refined particle dataset, is depicted in Fig. 5(d).
IV-B SOI Dynamics Model
@@
To characterize the dynamics of SOI particles, as illustrated in Fig. 4(b), we begin to construct a particle graph and introduce the graph neural networks to model the SOI dynamics. To construct a particle graph, we denote a graph representing SOI observations as $\mathcal{G}(\mathbf{\hat{o}}_{t})=(V_{t},E_{t})$, and the graph vertices $V_{t}$ correspond to the SOI’s particles $\mathbf{p}_{i,t}$. Each particle is expressed as $\mathbf{p}_{i,t}=\langle\mathbf{x}_{i,t},\mathbf{b}_{i,t}\rangle$, where the position and attributes of the particle $i$ at time $t$ are denoted as $\mathbf{x}_{i,t}$ and $\mathbf{b}_{i,t}$, respectively. Edges $E_{t}$ link vertices dynamically, based on spatial relationships, connecting all neighboring particles within a predefined range. Edge relations are captured by $\mathbf{e}_{j}=\langle m_{j},n_{j},\mathbf{c}_{j}\rangle$, with $1\leq m_{j},n_{j}\leq|\mathbf{P}_{t}|$ denoting the indices of the connected particles, $j$ being the edge index, and $\mathbf{c}_{j}$ describing the type of connection, be it internal structural or handle-to-rim linkages.
@@
The goal of introducing GNNs is to simulate the dynamics of the SOI and to predict subsequent states from a short historical sequence of SOI particle graphs, formalized as:
$\mathcal{G}(\mathbf{\hat{o}}_{t+1})=f_{\theta_{\text{dyn}}}\left(\mathcal{G}(\mathbf{\hat{o}}_{t-h:t}),\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{t}\right)$ (2)
@@
$\phi^{\mathbf{p}},\psi^{\mathbf{p}}$ $=\arg\min_{\phi^{\mathbf{v}},\psi^{\mathbf{v}}}\operatorname{Dist}[V-(\phi^{\mathbf{p}}\circ\psi^{\mathbf{p}})(P)]$ (3)
$\phi^{\mathbf{e}},\psi^{\mathbf{e}}$ $=\arg\min_{\phi^{\mathbf{e}},\psi^{\mathbf{e}}}\operatorname{Dist}[E-(\phi^{\mathbf{e}}\circ\psi^{\mathbf{e}})(E)]$
@@
$\mathbf{\hat{e}}_{k,t}$ $=\psi^{\mathbf{e}}(z_{j,t}^{\mathbf{e}})_{j=1,\cdots,|E_{t}|}$ (5)
$\hat{\mathbf{p}}_{i,t+1}$ $=\psi^{\mathbf{p}}\left(z_{i,t}^{\mathbf{p}},\sum_{j\in\mathcal{N}_{i}}\mathbf{\hat{e}}_{k,t}\right)_{i=1,\cdots,|\mathbf{P}_{t}|}$
@@
Our training data, derived from point clouds, lacks a consistent point-to-point mapping across each frame, thus precluding the use of particle-wise loss functions. To quantify the resemblance between two sets of SOI particle distributions, we investigate two distinct loss functions. The first is the widely-adopted Chamfer distance (CD), computed between two particle sets $\mathbf{P}_{1},\mathbf{P}_{2}\subseteq\mathbb{R}^{3}$ as follows:
@@
$\small\mathcal{L}_{\text{EMD}}(\mathbf{P}_{1},\mathbf{P}_{2})=\min_{\phi:\mathbf{P}_{1}\rightarrow\mathbf{P}_{2}}\sum_{\mathbf{x}_{1}\in\mathbf{P}_{1}}\|\mathbf{x}_{1}-\phi(\mathbf{x}_{1})\|_{2}$ (7)
where $\phi:\mathbf{P}_{1}\rightarrow\mathbf{P}_{2}$ denotes a bijective mapping. The EMD represents a solution to the assignment problem, ensuring a unique and stable optimal bijection $\phi$ for almost every pair of particle sets, invariant to infinitesimally small point displacements. In essence, EMD in our framework aligns distributions while mitigating point cloud anomalies through the definition of bijective correspondence. Regarding the Chamfer distance, it’s worth noting that its use here is somewhat liberal, as it does not fulfill the triangle inequality property. Our composite loss function integrates these distances in a weighted fashion: $\mathcal{L}(\mathbf{P}_{1},\mathbf{P}_{2})=\alpha\mathcal{L}_{\text{CD}}(\mathbf{P}_{1},\mathbf{P}_{2})+\beta\mathcal{L}_{\text{EMD}}(\mathbf{P}_{1},\mathbf{P}_{2})$. Empirical evaluations suggest that the optimal weights are $\alpha=0.85$ and $\beta=0.15$.
@@ IV-C Model Predictive Control
Upon training our SOI-centric latent dynamics model, we integrate a model predictive control (MPC) approach to control the robotic gripper in manipulating the fabric bag, as depicted in Fig. 4(c). We simplify the gripper’s action space into a parameterized form: $(x,y,z,r_{z})$, with $\{x,y,z\}$ representing the end-effector’s position interacting with the bag handles, and $r_{z}$ indicating the gripper’s rotation around the vertical $z$ axis. We omit $r_{x}$ and $r_{y}$ rotations based on empirical evidence suggesting minimal SOI deformation from these movements. A goal-oriented MPC is employed as below, using $\mathbf{g}$ to symbolize the desired SOI shape and $\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{0:{T-1}}$ to represent the sequence of action pairs derived via MPC, with $T$ as the planning horizon.
$\min_{\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{0:T-1}}$ $\mathcal{J}\left(\mathbf{\hat{o}}_{T},\mathbf{g}\right)=\min_{\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{0:T-1}}\mathcal{L}\left(\mathbf{\hat{P}}_{T},\mathbf{P}_{\mathbf{g}}\right)$ (8)
s.t. $\mathcal{G}(\mathbf{\hat{o}}_{t+1})=f_{\theta_{\text{dym}}}\left(\mathcal{G}(\mathbf{\hat{o}}_{t-h:t}),\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{t}\right)$
$\mathbf{P}_{\mathbf{g}}=\mathcal{G}(\mathbf{g})$
$\mathbf{\hat{P}}_{t}=\mathcal{G}(\mathbf{\hat{o}}_{t})$
$t=0,1,\ldots,T-1$
The objective is to identify action sequences that minimize the hybrid distance to the goal, expressed as $\mathcal{J}(\mathbf{\hat{g}}_{T},\mathbf{g})\equiv\mathcal{L}(\mathbf{\hat{P}}_{T},\mathbf{P}_{\mathbf{g}})$. Gradient-based trajectory optimization is utilized to ascertain the minimal-cost trajectory, showcased in Fig. 4(c). Initial random shooting in the simplified action space is followed by cost calculation via the GNN-driven dynamics model. Subsequently, gradient refinement with the limited-memory BFGS method [32] is conducted on the most cost-effective trajectories, employing the same loss function used during the dynamics model’s training phase.

### v2

@@ IV-A Global Particle Sampling from Raw Observation
To address the challenge of real-time extracting meaningful particle representations of the SOI from significantly occluded visual data, we introduce a comprehensive global particle sampling approach, which is illustrated in Fig. 5.
Preprocessing: The initial phase of our methodology involves preprocessing of the raw point cloud data obtained from RealSense D435 cameras. With appropriate calibration, we can transform the raw RGB-D images into point cloud data (PCD) (see Fig. 5(a)) with the cameras’ intrinsic and extrinsic parameters. Then, we focus on the SOI (i.e., opening rim) of the fabric bag and employ a color-based segmentation algorithm by performing a specific color filtering within the HSV color space to isolate the rim from the rest of the bag (see Fig. 5(b)). To enhance the processing efficiency in the following steps, we further reduce the points to satisfy the real-time performance using a parallel Voxel Grid Downsampling (VGD) by dividing the point cloud into a regular grid of 3D voxels (voxel size denoted by $v_{s}$) and replaces all the points within each voxel with their centroid or average. Then, followed by a GPU-accelerated implementation of the Statistical Outlier Removal (SOR) with a $k$-nearest neighbors to remove the noise and further improve the data quality (see Fig. 5(c)). In general, this step is intended to significantly reduce the size and complexity of the point cloud to an appropriate scale (typically, around 500 points) for subsequent surface reconstruction.
Surface Reconstruction and Refinement: To reconstruct the surface of SOI from its point cloud in a high efficiency, we leverage ball pivoting algorithm that is a preferred technique for its efficiency in managing point clouds. To achieve real-time surface reconstruction, we parallelize the algorithm, which allows the BPA to utilize multiple cores or processors simultaneously, thereby accelerating the reconstruction process. Specifically, we employ an Octree-based data structure to partition the point cloud into smaller subsets, each of which can be processed independently. For example, we can compute point normal (see Fig. 5(d)) via Octree-based nearest-neighbor search methods in each partitioned subset. Then we refine the reconstructed surface based on topological priors to get rid of non-manifold edges and vertices which can be easily detected by Open3D [^1] https://www.open3d.org/docs/release/tutorial/geometry/mesh.html. Non-manifold edges and vertices in a mesh are those that do not comply with the manifold criteria, which typically require that each edge belong to exactly two faces and each vertex to form a closed fan of faces around it. By removing non-manifold edges and vertices, we can ensure a well-defined volume to further improve the following resampling accuracy.
@@
Resampling: In the final stage of our process, our objective is to generate a high-quality particle set that accurately represents the geometry of the SOI-based on the reconstructed mesh. To achieve this, we employ uniform sampling techniques to selectively reduce the SOI point cloud in to an appropriate fixed scale, maintaining a uniform distribution of the generated particles. This ensures a manageable quantity conducive to Graph Neural Network (GNN) training. The computation of this process, resulting in a refined particle dataset, is depicted in Fig. 5(f).
IV-B SOI-based Neural Dynamics Model
@@
To characterize the dynamics of SOI particles, as illustrated in Fig. 4(b), we begin to construct a particle graph and introduce the graph neural networks to model the SOI dynamics. To construct a particle graph, we denote a graph based the SOI observation to represent the SOI state as $\mathbf{y}_{t}=\mathcal{G}(\mathbf{o}^{soi}_{t})=(\mathbf{P}_{t},\mathbf{E}_{t})$, and the graph vertex set $\mathbf{P}_{t}$ correspond to the SOI’s particles $\mathbf{p}_{i,t}$. Each particle is expressed as $\mathbf{p}_{i,t}=\langle\mathbf{x}_{i,t},\mathbf{b}_{i,t}\rangle$, where the position and attributes of the particle $i$ at time $t$ are denoted as $\mathbf{x}_{i,t}$ and $\mathbf{b}_{i,t}$, respectively. Attributes of the particles $\mathbf{b}_{i,t}$ are classified into handle particles and non-handle particles. The handle particles are identified by a open ball function as below:
$\mathcal{B}_{r}(\mathbf{x}_{handle})=\{\mathbf{p}_{i}\in\mathbf{P}~{}|~{}d(\mathbf{p}_{i},\mathbf{x}_{handle})<r\}$ (2)
where $d(\mathbf{p}_{i},\mathbf{x}_{handle})$ represents the distance between SOI particles and bag handle position in Euclidean space. The handle position is also identified by HSV filtering in color space. Introducing this particle attribute significantly enhances the SOI neural dynamics model. The neural dynamics model relies on understanding how different parts (nodes) of the SOI, which lie on the deformable object, behave under robot manipulations. By identifying the position of the handle, which has a strong relationship with the enforced robot actions, the model can accurately locate and classify these critical nodes. Therefore, node classification allows the model to learn and predict the dynamics of specific particles of the SOI, particularly those near the handle, with greater accuracy. Edges $\mathbf{E}_{t}$ link vertices dynamically, based on spatial relationships, connecting all neighboring particles within a predefined range. Edge relations are captured by $\mathbf{e}_{j}=\langle m_{j},n_{j},\mathbf{c}_{j}\rangle$, with $1\leq m_{j},n_{j}\leq|\mathbf{P}_{t}|$ denoting the indices of the connected particles, $j$ being the edge index, and $\mathbf{c}_{j}$ describing the type of connection, be it internal structural or handle-to-rim linkages.
The goal of constructing a forward dynamics model is to predict SOI subsequent state $\mathbf{y}_{t+1}$ based on a short historical sequence of SOI states, denoted by:
$\mathbf{y}_{t+1}=f_{\theta_{\text{dyn}}}\left(\mathbf{y}_{t-h:t},\mathbf{a}_{t}\right)$ (3)
@@
With the application of particle graph to represent the SOI states, we can introduce GNNs to simulate the neural dynamics of the SOI to predict subsequent states from a short historical sequence of SOI particle graphs, extended as:
$\mathcal{G}(\mathbf{o}^{soi}_{t+1})=f_{\theta_{\text{dyn}}}\left(\mathcal{G}(\mathbf{o}^{soi}_{t-h:t}),\{\mathbf{a}^{0},~{}\mathbf{a}^{1}\}_{t}\right)$ (4)
@@
$\phi^{\mathbf{p}},\psi^{\mathbf{p}}$ $=\arg\min_{\phi^{\mathbf{p}},\psi^{\mathbf{p}}}\operatorname{Dist}[\mathbf{p}_{i}-(\phi^{\mathbf{p}}\circ\psi^{\mathbf{p}})(\mathbf{p}_{i})]$ (5)
$\phi^{\mathbf{e}},\psi^{\mathbf{e}}$ $=\arg\min_{\phi^{\mathbf{e}},\psi^{\mathbf{e}}}\operatorname{Dist}[\mathbf{e}_{j}-(\phi^{\mathbf{e}}\circ\psi^{\mathbf{e}})(\mathbf{e}_{j})]$
@@
$\mathbf{\hat{e}}_{j,t}$ $=\psi^{\mathbf{e}}(z_{j,t}^{\mathbf{e}})_{j=1,\cdots,|\mathbf{E}_{t}|}$ (7)
$\hat{\mathbf{p}}_{i,t+1}$ $=\psi^{\mathbf{p}}\left(z_{i,t}^{\mathbf{p}},\sum_{j\in\mathcal{N}_{i}}\mathbf{\hat{e}}_{j,t}\right)_{i=1,\cdots,|\mathbf{P}_{t}|}$
@@
Since our training data is based on the particle sets generated from raw point cloud data with our proposed global particle sampling process, it lacks a consistent point-to-point mapping across each frame. Therefore, to quantify the similarity between two sets of SOI particle distributions, we investigate two permutation-invariant loss functions. The first is the widely adopted Chamfer distance (CD), computed between two particle sets $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ as follows:
@@
$\small\mathcal{L}_{\text{EMD}}(\mathbf{P}_{1},\mathbf{P}_{2})=\min_{\lambda:\mathbf{P}_{1}\rightarrow\mathbf{P}_{2}}\sum_{\mathbf{x}_{1}\in\mathbf{P}_{1}}\|\mathbf{x}_{1}-\lambda(\mathbf{x}_{1})\|_{2}$ (9)
where $\lambda:\mathbf{P}_{1}\rightarrow\mathbf{P}_{2}$ denotes a bijective mapping. The EMD represents a solution to the assignment problem, ensuring a unique and stable optimal bijection $\lambda$ for almost every pair of particle sets, invariant to infinitesimally small point displacements. Note that the bijection $\lambda$ will be computed between two particle set for every time which means the bijection will be regenerated. To ensure real-time performance, we will limit the cardinality of the particle set by configuring the GSP to generate a small number of particles. This reduction in particle count helps minimize the computational burden, allowing us to meet the necessary requirements for a real-time controller. In essence, EMD in our framework aligns distributions while mitigating point cloud anomalies through the definition of bijective correspondence. Regarding the Chamfer distance, it’s worth noting that its use here is somewhat liberal, as it does not fulfil the triangle inequality property. Our composite loss function integrates these distances in a weighted fashion: $\mathcal{L}(\mathbf{P}_{1},\mathbf{P}_{2})=\alpha\mathcal{L}_{\text{CD}}(\mathbf{P}_{1},\mathbf{P}_{2})+\beta\mathcal{L}_{\text{EMD}}(\mathbf{P}_{1},\mathbf{P}_{2})$. Empirical evaluations suggest that the optimal weights are $\alpha=0.15$ and $\beta=0.85$.
@@ IV-C Model Predictive Control
Figure 6: The MPC framework integrates our trained SOI-based neural dynamics model for bimanual deformable bag manipulation.
@@
Upon training our SOI-centric neural dynamics model, we integrate a model predictive control (MPC) approach to control the robotic gripper in manipulating the fabric bag, as depicted in Fig. 10. We simplify the gripper’s action space into a parameterized form: $(x,y,z,r_{z})$, with $\{x,y,z\}$ representing the end-effector’s position interacting with the bag handles, and $r_{z}$ indicating the gripper’s rotation around the vertical $z$ axis. We considered only $r_{z}$ in our action space based on our observation that $r_{z}$ induces larger SOI configuration changes compared to $r_{x}$ and $r_{y}$. A goal-oriented MPC with $T$ as the planning horizon for the desired SOI observation $\mathbf{o}^{soi}_{*}$, denoted as the particle set $\mathbf{P}_{\mathbf{o}^{soi}_{*}}$, is employed as below:
$\small\min_{\{\mathbf{a}^{0},~{}\mathbf{a}^{1}\}_{0:T-1}}$ $\sum_{t=0}^{T}\mathcal{L}\left(\mathbf{P}_{\mathbf{o}^{soi}_{t}},\mathbf{P}_{\mathbf{o}^{soi}_{*}}\right)$ (10)
s.t. $\mathcal{G}(\mathbf{o}^{soi}_{t+1})=f_{\theta_{\text{dyn}}}\left(\mathcal{G}(\mathbf{o}^{soi}_{t-h:t}),\{\mathbf{a}^{0},~{}\mathbf{a}^{1}\}_{t}\right)$
$d_{lb}\leq||\mathbf{R}^{0}_{t}-\mathbf{R}^{1}_{t}||\leq d_{ub}$
$||\langle\mathbf{a}^{0},~{}\mathbf{a}^{1}\rangle_{t}||\leq\epsilon$
where we constrain the distance between left robot end-effector pose $\mathcal{R}^{0}$ and right robot end-effector pose $\mathcal{R}^{1}$ to fall within a specified range from the lower bound $d_{lb}$ to the upper bound $d_{ub}$ so as to prevent potential collisions between the manipulators and safeguard the fabric from tearing due to excessive force. Additionally, we ensure that the magnitude of the robot’s actions does not exceed $\epsilon$ to enforce a smooth and continuous trajectory.
As illustrated in Fig. 4(c), gradient-based trajectory optimization is utilized to identify the trajectory with the lowest defined cost. First, random shooting within the simplified actions is performed, followed by the computation of costs using the SOI-based latent dynamics model. Subsequently, we employ the limited-memory BFGS method [28] on the trajectories with the lowest cost to optimize the manipulation actions with gradients, using the same loss function during the dynamics model’s training process. The global convergence is guaranteed as proved in [29] and we empirically demonstrated the effectiveness through experiments.
V Experiments
V-A Experiment Setup
As shown in Fig. 7, we present the general experimental setup used to validate our SOI-based dynamics model for bimanual deformable bag manipulation tasks. The setup includes four RealSense D435 RGB-D cameras, each positioned at a corner, to capture RGB-D images of the deformable fabric bag from multiple angles at a frequency of 30Hz and a resolution of 640$\times$480. Two Dobot CR5 robotic manipulators, equipped with 7 degrees of freedom, grasp different SOIs using 3D-printed grippers secured with zip ties or binder clips. To validate the effectiveness of our proposed approach, three SOIs of the deformable objects, including the opening rim of a bag, the bottom hem and the collar of a T-shirt, have been modified for better perception: it is cut and sewn with the same type of fabric but in contrasting colors, which facilitates the detection while maintaining the consistency of the SOI’s dynamic behavior. Additionally, a top-down camera and a front-facing camera are employed to capture the experimental process from various perspectives. To support real-time point cloud process, particle sampling and MPC optimization, we utilize a workstation with Ubuntu 20.04, an Intel Core i9-14900K CPU (24 cores, 24 threads), an Nvidia RTX 4090 GPU, 128GB of DDR4/DDR5 RAM, and a 2TB NVMe SSD for storage. Table I presents the detailed computation time for the proposed global particle sampling and GNN-based neural dynamics model-integrated MPC optimization process. The total computation time required to generate a robot action command is approximately 63.75ms, theoretically allowing the robot to execute up to 15.69 action commands per second to manipulate deformable objects. However, to maintain consistency with the action execution frame rate used during the training of our GNN-based neural dynamics model, we standardize the execution frame rate of robot action commands to 10 FPS.
TABLE I: Computation time of the proposed global particle sampling and MPC optimization process.
Process Computation Time (ms) $\downarrow$
Point Cloud Registration $3.13$ $\mathrm{p}\mathrm{m}$  0.26
SOI Extraction $6.94$ $\mathrm{p}\mathrm{m}$  0.47
Down-sampling $7.05$ $\mathrm{p}\mathrm{m}$  0.73
Outlier Removal $8.75$ $\mathrm{p}\mathrm{m}$  0.61
Ball Pivoting $15.65$ $\mathrm{p}\mathrm{m}$  1.42
Particle Sampling $6.08$ $\mathrm{p}\mathrm{m}$  0.37
MPC Computation $17.12$ $\mathrm{p}\mathrm{m}$  1.79
GNN Inference $0.37$ $\mathrm{p}\mathrm{m}$  0.03